OpenStudy (anonymous):

PLEASE HELP: Solve x - 2y = 6 3x - 6y = 18 Question 8 options: 1) (2, -2) 2) (3, -3/2) 3) No Solutions 4) Infinitely Many Solutions

jimthompson5910 (jim_thompson5910):

In equation 1, solve for x to get x = 2y+6 3x-6y = 18 3(2y+6)-6y = 18 6y+18-6y = 18 18 = 18 Since this equation is ALWAYS true, this means that there are an infinite number of solutions. So the answer is choice 4)